OpenStudy (anonymous):

What is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)?

OpenStudy (anonymous):

Are you trying to find the circumference or radius?

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2 = r^2 is the standard equation for a circle for starters. (h,k) is the center of the circle; the point given was (-2,1) -2 is point h, and 1 is point k. R^2 is the radius squared. H corresponds with X and K corresponds with K for a side note. So using the points (-2,1) and (-4,1) to find it's radius, the differences between -2 and -4 is 2 points. So 2 is the radius of the circle. So now we got the points (-2,1) and the radius of the circle. Now all we have to do is put it in equation form. (-2, 1) r=2 would become (x-h)^2 + (y-k)^2 = r^2. In the end, it is (x+2)^2 + (y-1)^2=2^2